### v2

# Split Matching for Inductive Zero-shot Semantic Segmentation

Abstract
Zero-shot Semantic Segmentation (ZSS) aims to segment categories that are not annotated during training. While fine-tuning vision-language models has achieved promising results, these models often overfit to seen categories due to the lack of supervision for unseen classes. As an alternative to fully supervised approaches, query-based segmentation has shown great latent in ZSS, as it enables object localization without relying on explicit labels. However, conventional Hungarian matching, a core component in query-based frameworks, needs full supervision and often misclassifies unseen categories as background in the setting of ZSS. To address this issue, we propose Split Matching (SM), a novel assignment strategy that decouples Hungarian matching into two components: one for seen classes in annotated regions and another for latent classes in unannotated regions (referred to as unseen candidates). Specifically, we partition the queries into seen and candidate groups, enabling each to be optimized independently according to its available supervision. To discover unseen candidates, we cluster CLIP dense features to generate pseudo masks and extract region-level embeddings using CLS tokens. Matching is then conducted separately for the two groups based on both class-level similarity and mask-level consistency. Additionally, we introduce a Multi-scale Feature Enhancement (MFE) module that refines decoder features through residual multi-scale aggregation, improving the model’s ability to capture spatial details across resolutions. SM is the first to introduce decoupled Hungarian matching under the inductive ZSS setting, and achieves state-of-the-art performance on two standard benchmarks.
1 Introduction
Semantic segmentation [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2024b)Chen, Deguchi, Zhang, Zheng, and Murase] serves as a fundamental task for fine-grained scene understanding in various applications, such as autonomous driving, robotic perception, and medical imaging [Zheng et al.(2022)Zheng, Fu, Xie, Chen, Wang, and Sham, Chen et al.(2022)Chen, Fu, Xie, Zheng, Geng, and Sham, Zheng et al.(2024a)Zheng, Luo, Fu, Liu, and Wang, Xie et al.(2023)Xie, Fu, Zheng, Zheng, Sham, and Wang]. Existing approaches can be broadly classified into feature-based and query-based methods. Feature-based methods treat semantic segmentation as a per-pixel classification problem, where dense features extracted from the backbone are directly finetuned for pixel-wise label prediction. In contrast, query-based segmentors employ a set of discrete, learnable vectors, referred to as queries, to jointly predict class labels and class-agnostic masks. These queries are passed through a transformer decoder to produce class scores and simultaneously interact with dense backbone features to generate the corresponding masks.
Figure 1: Comparisons between existing query-based segmentation models that fail to match the unseen candidates and the proposed split matching.
[…]
To address the limitations of existing approaches, particularly their tendency to overfit to seen classes, we explore query-based segmentation as a more effective alternative for zero-shot settings. Unlike feature-based methods that classify each pixel independently, query-based methods treat each object as a learnable query and perform mask-level classification, enabling better modeling of object-level semantics and localization of latent classes without requiring explicit labels. Despite this advantage, current query-based methods struggle in ZSS scenarios, where the absence of annotations for unseen classes during training causes queries to be predominantly assigned to seen categories. As a result, queries corresponding to unseen objects remain unmatched, are treated as background, and thus fail to detect unseen categories. As illustrated in the left of Fig. 1, this matching imbalance introduces a strong bias toward seen classes, leaving the model unable to train or preserve representations for unseen classes, limiting the full latent of query-based segmentation for ZSS.
To address the challenge of segmenting unseen classes, we propose Split Matching (SM) for query-based segmentation. Specifically, we divide the queries into two groups: seen queries, which segment annotated seen classes, and candidate queries, which target latent classes in unannotated regions (referred to as unseen candidates). We first apply multi-scale K-Means clustering [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] on CLIP dense features to generate pseudo masks that localize unseen candidates. The corresponding image patches are then cropped and fed into the frozen CLIP visual encoder to obtain CLS tokens, which are subsequently fused with semantic embeddings to form joint vision-language representations. We compute class similarity between the joint embeddings and query predictions, and measure mask similarity using the pseudo masks. Hungarian matching is then applied separately to the seen and candidate queries based on the combined similarities. To further enhance feature quality, we introduce a Multi-scale Feature Enhancement (MFE) module, which refines the transformer decoder’s key and value via residual multi-scale aggregation.
Different from existing methods that rely on dense features extracted from CLIP [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase] and struggle to optimize unannotated regions due to the absence of explicit supervision, our method introduces Split Matching (SM), which separates queries into seen and candidate groups, enabling targeted label assignment even in unannotated areas. This design effectively mitigates the common issue of misclassifying unseen objects as background—a key limitation of prior methods. Moreover, unlike open-vocabulary segmentation approaches [Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai] that perform Hungarian matching with fully annotated data, our method operates entirely under the zero-shot setting, without requiring any pixel-level labels for unseen classes. To the best of our knowledge, we are the first to explicitly separate seen and candidate queries via Hungarian matching under the inductive ZSS. To summarize, our contributions are:
- We propose Split Matching (SM), the first query-based assignment framework for zero-shot segmentation that separates seen and unseen candidates using pseudo masks derived from CLIP features without requiring fully labeled datasets.
- We introduce a Multi-scale Feature Enhancement (MFE) module that enriches decoder features via residual fusion, improving query interaction and segmentation quality.
- Our method achieves state-of-the-art performance on PASCAL VOC and COCO-Stuff, demonstrating strong generalization to unseen classes and balanced performance.
2 Related Works
Semantic Segmentation assigns a class label to each pixel in an image [Zheng et al.(2023b)Zheng, Zhu, Liu, Cao, Fu, and Wang, Zheng et al.(2023a)Zheng, Pan, Luo, and Wang, Zheng et al.(2024e)Zheng, Zhou, Vasilakos, and Wang, Zheng et al.(2024b)Zheng, Lyu, and Wang, Zheng et al.(2024d)Zheng, Zhou, Vasilakos, and Wang, Zheng et al.(2024c)Zheng, Lyu, Zhou, and Wang, Zhu et al.(2024)Zhu, Xiao, Shi, Xu, and Zheng, Liao et al.(2025b)Liao, Zheng, Lyu, Xue, Cao, Wang, Yang, and Hu, Liao et al.(2025a)Liao, Lei, Zheng, Moon, Wang, Wang, Paudel, Gool, and Hu, Zhong et al.(2025)Zhong, Zheng, Liao, Lyu, Chen, Wu, Zhang, and Hu]. Traditional CNN-based methods [Long et al.(2015)Long, Shelhamer, and Darrell, Chen et al.(2017)Chen, Papandreou, Kokkinos, Murphy, and Yuille, Chen et al.(2018)Chen, Zhu, Papandreou, Schroff, and Adam] enhance per-pixel classification through dilated convolutions and multi-scale context aggregation but struggle to capture long-range dependencies. With the rise of Vision Transformers [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.], encoder-based models [Xie et al.(2021)Xie, Wang, Yu, Anandkumar, Alvarez, and Luo, Guo et al.(2022)Guo, Lu, Hou, Liu, Cheng, and Hu] have shown strong capabilities in modeling global context. Inspired by DETR [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko], query-based frameworks such as MaskFormer [Cheng et al.(2021a)Cheng, Schwing, and Kirillov] and Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] reformulate segmentation as set prediction using object queries and Hungarian matching. Although these models achieve impressive results under full supervision, they are not directly applicable to zero-shot segmentation (ZSS) due to their reliance on complete annotations and inherent bias toward seen classes. To address this, we propose Split Matching, the first method to explicitly separate seen and candidate queries under the inductive ZSS setting. By decoupling the matching process for annotated and unannotated regions, our approach enables targeted supervision and significantly improves generalization to unseen categories, even in the absence of pixel-level annotations.
Zero-shot and Open-Vocabulary Segmentation. Although semantic segmentation has made remarkable progress, it still relies heavily on large-scale pixel-level annotations, which are expensive and time-consuming to obtain. To alleviate this, two related directions have emerged: zero-shot segmentation (ZSS) [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase, Chen et al.(2024a)Chen, Deguchi, Zhang, and Murase, Ye et al.(2024)Ye, Oh, et al.] and open-vocabulary segmentation (OVS) [Xu et al.(2023)Xu, Zhang, Wei, Hu, and Bai, Yu et al.(2023)Yu, He, Deng, Shen, and Chen, Wu et al.(2024)Wu, Zhang, Xu, Jin, Li, Liu, and Loy, Zhou et al.(2022)Zhou, Loy, and Dai]. Both paradigms aim to improve generalization by leveraging large-scale vision-language models such as CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], either by introducing lightweight adapters [Houlsby et al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De Laroussilhe, Gesmundo, Attariyan, and Gelly] or designing task-specific visual prompts [Jia et al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim]. Despite this shared goal, their setups differ fundamentally. ZSS is defined under a partially labeled training regime, where only a subset of classes is annotated and the rest are marked as ignored. Evaluation is conducted on the same domain, including both seen and unseen categories. OVS, on the other hand, assumes access to fully labeled training data and evaluates on datasets with novel categories and distribution shifts. While OVS benefits from full supervision and can leverage existing segmentation architectures, ZSS faces the challenge of incomplete supervision. Unannotated regions corresponding to unseen classes are often treated as background during training, making it difficult for models to learn discriminative representations for unseen concepts. To tackle this issue, we propose Split Matching (SM), a label assignment strategy tailored for query-based models in the zero-shot setting. SM dynamically aligns predicted queries with pseudo labels derived from external vision-language features, enabling the discovery of unseen objects even in the absence of ground-truth annotations.
Figure 2: Overview of the proposed method. During training, input images are fed into a trainable backbone to extract dense features, while a frozen CLIP encoder provides CLIP features. These are used to generate class embeddings $\textbf{C}_{u}$ and pseudo masks $\textbf{Y}_{p}$ for both seen and unseen classes. The dense features serve as keys and values in a transformer decoder, which interacts with queries Q to predict masks and query features. These are aligned with pseudo labels via the Split Matching. A Multi-scale Feature Enhancement module further refines features with auxiliary loss $\mathcal{L}_{MFE}$.
3 Methods
3.1 Preliminaries
Task Definition. ZSS aims to segment both seen and unseen classes without pixel-level annotations for unseen classes after training. Formally, let $\mathcal{D}=\{\textbf{X}^{i},\textbf{Y}_{s}^{i}\}_{i=1}^{M}$ represent a dataset of images X and their pixel-level annotations $\textbf{Y}_{s}$ for seen classes. Additionally, let $\textbf{A}\in\mathbb{R}^{N\times C}$ denote the semantic (text) embeddings of all classes, divided into seen $\textbf{A}_{s}\in\mathbb{R}^{N_{s}\times C}$ and unseen $\textbf{A}_{u}\in\mathbb{R}^{N_{u}\times C}$, such that $(\textbf{A}_{s}\cap\textbf{A}_{u}=\varnothing)$ and $N_{s}+N_{u}=N$ where $C$ indicates the channel number and $N$ indicates the number of classes in the dataset. Our method applies the Inductive settings, where unseen embeddings $\textbf{A}_{u}$ are inaccessible during training and evaluates model performance on seen and unseen classes during inference. Meanwhile, all training images are preserved during training, and regions corresponding to unseen classes are consistently labeled as “ignored”, ensuring that no unseen information is used.
Method Overview. Our core idea is to mitigate the seen-class bias caused by incomplete annotations by decoupling the optimization of seen and unannotated regions, thereby enabling better discovery of unseen classes without degrading seen-class performance. As illustrated in Fig. 2, we divide the query space into two parts: seen queries and candidate queries, which are responsible for segmenting seen classes in the annotated areas and unseen candidates in the unannotated areas, respectively. Given an input image, a trainable backbone extracts dense features that serve as keys and values for a transformer decoder. A set of randomly initialized queries $\mathbf{Q}$ interacts with these features to produce segmentation outputs, which are then split into $\mathbf{Q}_{s}$ (seen queries) and $\mathbf{Q}_{u}$ (candidate queries). To guide $\mathbf{Q}_{u}$ towards discovering unseen candidates, the input image is also passed through a frozen CLIP encoder. From the CLIP features, we generate pseudo masks $\mathbf{Y}_{u}$ and corresponding class embeddings $\mathbf{C}_{u}$ by extracting the CLS tokens of the unseen candidates. We then introduce Split Matching (Sec.3.2), which assigns labels to both seen and candidate queries by computing similarity between model predictions, pseudo masks, and class embeddings. To further enhance feature quality, we propose a Multi-scale Feature Enhancement module (Sec.3.3), which refines the dense features extracted by the backbone.
Although our method utilizes features from unannotated regions, it strictly adheres to the inductive zero-shot setting. At the beginning of training, all unannotated regions are uniformly treated as ‘ignored’ without introducing any class-specific supervision or bias. Unseen candidates are identified in a fully self-supervised manner, without any assumptions regarding the number, identity, or distribution of unseen classes. As a result, the model remains completely agnostic to unseen classes throughout the entire training process, ensuring that no unseen-related information is leaked.
3.2 Split Matching (SM)
To apply Hungarian matching in semantic segmentation, we need to split the ground-truth into class labels and their corresponding class-agnostic masks. We then compute the similarity between the predicted classes and masks and the ground-truth annotations. Based on these similarities, each ground-truth instance is assigned to a query and its associated mask. However, in ZSS, unannotated regions lack both ground-truth masks and class labels, making direct assignment infeasible. To address this issue, SM generates pseudo masks for latent classes (unseen candidates) in the unannotated areas and assigns them to candidate queries, while optimizing seen and candidate queries separately.
Figure 3: Overview of SM.
Figure 4: Overview of MFE.
To generate the pseudo masks and their corresponding class embeddings for unseen candidates, we use the multi-scale K-means and mask fusion methods from CLIP-ZSS [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase]. Specifically, given an image X, we feed it into a frozen CLIP visual encoder to obtain the CLIP dense features O. Then, we compute the seed for K-means methods with,
$\textbf{G}=\left\{\left.\sum_{u=i}^{i+s-1}\sum_{v=j}^{j+s-1}\frac{\textbf{O}[u,v]}{s^{2}}\right\rvert\,i\in I,j\in J\right\},$ (1)
where $I=\{0,[s/2],\dots,[H-s]\}$ and $J=\{0,[s/2],\dots,[W-s]\}$ are index sets, []delimited-[][$\cdot$][ ⋅ ] denotes rounding, and $s\in S$ is the window size. After we obtain G, we merge the masks corresponding to G by the mask fusion algorithm [Chen et al.(2023)Chen, Deguchi, Zhang, Zheng, and Murase], which merge the semantically similar masks by the cosine similarity among cluster centroids, to obtain pseudo label $\textbf{Y}_{u}\in[0,1]^{U\times H\times W}$ for unseen candidates where $U$ indicates the number of unseen candidates. After obtaining $\textbf{Y}_{u}$, we use $\textbf{Y}_{u}$ to mask the input images X into $\textbf{X}_{m}\in\mathbb{R}^{U\times 3\times H\times W}$. Finally, we feed $\textbf{X}_{m}$ into the frozen CLIP visual encoder for $\textbf{C}_{u}\in\mathbb{R}^{U\times C}$ as class embeddings for unseen candidates.
After obtaining $\textbf{Y}_{u}$ and $\textbf{C}_{u}$, we apply split matching as illustrated in Fig. 4. Specifically, we first feed the randomly initialized queries Q into a transformer decoder to obtain the predictions for each query, denoted as $\textbf{P}=\{\textbf{P}_{s},\textbf{P}_{u}\}$. Here, we denote the predictions for seen and candidate queries as $\textbf{P}_{s}=\{\textbf{V}_{s},\textbf{M}_{s}\}$ and $\textbf{P}_{u}=\{\textbf{V}_{u},\textbf{M}_{u}\}$, respectively. Specifically, $\textbf{V}_{s}\in\mathbb{R}^{K_{s}\times C}$ and $\textbf{M}_{s}\in\mathbb{R}^{K_{s}\times H\times W}$ denote the features after transformer decoder and predicted masks for the $K_{s}$ seen queries. Similarly, $\textbf{V}_{u}\in\mathbb{R}^{K_{u}\times C}$ and $\textbf{M}_{u}\in\mathbb{R}^{K_{u}\times H\times W}$ correspond to the predictions for the $K_{u}$ candidate queries. In both cases, the queries are projected into a $C$-dimensional semantic space. Similar to the conventional hungarian matching, SM also consists of class matching and mask matching. For class matching, we concatenate the seen class embeddings $\textbf{A}_{s}$ (extracted from the CLIP textual encoder) and candidate-class embeddings $\textbf{C}_{u}$ to form a joint class embedding $\textbf{E}=\mathrm{cat}(\textbf{A}_{s},\textbf{C}_{u})\in\mathbb{R}^{(N_{s}+U)\times C}$, where ‘$\mathrm{cat}$’ denotes concatenation. Finally, we separately compute the similarity between the predictions and joint class embeddings for seen and candidate queries, respectively. $\textbf{S}_{s}=Sigmoid(\textbf{V}_{s}\cdot\textbf{E}^{\top}),\textbf{S}_{u}=Sigmoid(\textbf{V}_{u}\cdot\textbf{E}^{\top})$. Although both seen and candidate queries compute similarity against the unified embedding matrix $\mathbf{E}=\text{cat}(\mathbf{A}_{s},\mathbf{C}_{u})$, we apply Hungarian matching separately for seen and candidate queries. This allows each group to focus on relevant candidates while still benefiting from global semantic context.
[…]
$\sigma_{s}^{*}=\arg\min\sum_{i=0}^{K_{s}-1}\mathcal{L}_{\text{match}}(\textbf{S}_{s}^{\sigma(i)},\textbf{M}_{s}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{s}^{i}),\quad\sigma_{u}^{*}=\arg\min\sum_{i=0}^{K_{u}-1}\mathcal{L}_{\text{match}}(\textbf{S}_{u}^{\sigma(i)},\textbf{M}_{u}^{\sigma(i)},\textbf{E}_{i},\textbf{Y}_{u}^{i})$ (2)
After computing $\sigma_{s}^{*}$ and $\sigma_{u}^{*}$, we concatenate them into a unified assignment $\sigma^{*}$. The resulting assignment $\sigma^{*}$ is then used to optimize the matching loss across all queries. $\mathcal{L}_{match}$ is composed of two terms: a classification loss $\mathcal{L}_{cls}$ and a mask loss $\mathcal{L}_{mask}$:
$\mathcal{L}_{\text{match}}(\textbf{S},\textbf{M},\textbf{E},\textbf{Y})=\mathcal{L}_{cls}(\textbf{S}^{\sigma^{*}},\textbf{E})+\mathcal{L}_{mask}(\textbf{I}^{\sigma^{*}},\textbf{Y}),$ (3)
where $\textbf{S}\in\mathbb{R}^{K\times C}$ denotes classification scores over $C$ classes for $K$ queries, and $\textbf{M}\in\mathbb{R}^{K\times H\times W}$ represents the corresponding predicted masks. $\textbf{S}^{\sigma^{*}}$ and $\textbf{I}^{\sigma^{*}}$ are the classification scores and mask similarities after matching with the optimal assignment $\sigma^{*}$. To enhance the discriminability of candidate queries, we introduce a cosine similarity loss: $\mathcal{L}_{cos}=1-\mathrm{cos}(\textbf{V}_{u}^{\prime},\textbf{C}_{u})$, where $\textbf{V}_{u}^{\prime}$ are the matched candidate queries, and $\mathrm{cos}(a,b)$ denotes cosine similarity. The final training objective for split matching is $\mathcal{L}_{SM}=\mathcal{L}_{match}+\mathcal{L}_{cos}$.
3.3 Multi-scale Feature Enhancement
While SM effectively enables query-based approaches to tackle ZSS, the associated keys and values in transformer decoders used to locate corresponding regions in an image remain suboptimal, yet the current design lacks sufficient refinement, limiting the performance of the matching. To tackle this issue, we propose an Multi-scale Feature Enhancement (MFE) module, as illustrated in Fig. 4, designed to assist in identifying the relevant regions by effectively combining multi-scale features. The MFE leverages hierarchical features extracted by the pixel decoder to provide a comprehensive representation, capturing both fine-grained and global contextual information. The multi-scale outputs of the pixel decoder, $\textbf{F}=\left\{\textbf{F}_{0},\textbf{F}_{1},\textbf{F}_{2}\right\}$, represent features at finer resolutions, where $\textbf{F}_{0}$ has the coarsest resolution and $\textbf{F}_{2}$ the finest where $\textbf{F}_{i}\in\mathbb{R}^{C\times(H/r^{2-i})\times(W/r^{2-i})}$, with $r$ denoting the scale factor.
The fusion process begins by refining $\textbf{F}_{0}$, the lowest-resolution feature map, through a dense block consisting of a convolutional layer, group normalization [Wu and He(2018)], and a ReLU activation function. This block enhances local spatial relationships and prepares the feature map for integration with features at finer scales. The refined $\textbf{F}_{0}$ is then resized to align with the spatial dimensions of $\textbf{F}_{1}$, which undergoes a similar refinement process. The refined $\textbf{F}_{0}$ and $\textbf{F}_{1}$ are fused via element-wise addition, combining detailed lowest-resolution features with mid-level semantic information. This combined representation is resized again to match the spatial dimensions of $\textbf{F}_{2}$, the finest feature map in the hierarchy. Simultaneously, $\textbf{F}_{2}$ is processed through its own dense block to extract refined features. The previously fused representation of $\textbf{F}_{0}$ and $\textbf{F}_{1}$ is then combined with the refined $\textbf{F}_{2}$ to produce the final unified feature map, $\textbf{F}_{u}$. This hierarchical fusion ensures that fine-grained details from the highest-resolution features are effectively integrated with broader contextual cues. The final output, $\textbf{F}_{u}$, leverages complementary information across scales to provide a rich, unified representation that allows queries to accurately identify relevant regions in the image. After obtaining $\textbf{F}_{u}$, we optimize it through $\mathcal{L}_{MFE}=\mathcal{L}_{ce}(\textbf{F}_{u},\textbf{Y}_{p})+\mathcal{L}_{focal}(\textbf{F}_{u},\textbf{Y}_{p})$ where $\mathcal{L}_{ce}$ and $\mathcal{L}_{focal}$ are cross entropy and focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár].
3.4 Training and Inference
The total loss functions are $\mathcal{L}=\mathcal{L}_{SM}+\mathcal{L}_{MFE}$. During inference, we exclude the MFE module and utilize only the backbone and transformer decoder. Both trained seen and candidate queries, along with additional Randomly-initialized Queries (RQ) $\textbf{Q}_{r}$, are used to produce the final prediction. These RQs increase the density of queries in the feature space, enabling better coverage of unannotated regions and discovering additional latent candidates. The inference process is further illustrated in the Supplementary Materials.
4 Experiments
Dataset. To assess the effectiveness of our proposed method, we conduct experiments on the widely-used benchmark COCO-Stuff [Caesar et al.(2018)Caesar, Uijlings, and Ferrari] and PASCAL VOC [Everingham et al.(2015)Everingham, Eslami, Van Gool, Williams, Winn, and Zisserman], focusing on the task of zero-shot semantic segmentation (ZSS). We adopt the same seen and unseen category splits as in prior works [Ding et al.(2022)Ding, Xue, Xia, and Dai, Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu, Zhou et al.(2022)Zhou, Loy, and Dai]. Specifically, COCO-Stuff consists of a total of 171 classes with 156 seen and 15 unseen classes according to the standard protocol. The dataset includes 118,287 images for training and 5,000 images for testing. PASCAL VOC contains 10,582 images for training and 1,449 images for validation, including 15 seen and 5 unseen classes.
      Models       Backbone       PASCAL VOC       COCO-Stuff
      hIoU       sIoU       uIoU       hIoU       sIoU       uIoU
      SPNet [Xian et al.(2019)Xian, Choudhury, He, Schiele, and Akata]       ResNet101 [He et al.(2016)He, Zhang, Ren, and Sun]       26.1       78.0       15.6       14.0       35.2       8.7
      ZS3 [Bucher et al.(2019)Bucher, Vu, Cord, and Pérez]       28.7       77.3       17.7       15.0       34.7       9.5
      CaGNet [Gu et al.(2020)Gu, Zhou, Niu, Zhao, and Zhang]       39.7       78.4       26.6       18.2       33.5       12.2
      SIGN [Cheng et al.(2021b)Cheng, Nandi, Natarajan, and Abd-Almageed]       41.7       75.4       28.9       20.9       32.3       15.5
      Joint [Baek et al.(2021)Baek, Oh, and Ham]       45.9       77.7       32.5       -       -       -
      ZegFormer [Ding et al.(2022)Ding, Xue, Xia, and Dai]       73.3       86.4       63.6       34.8       36.6       33.2
      Zzseg [Xu et al.(2022)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai]       ViT-B [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.]       77.5       83.5       72.5       37.8       39.3       36.3
      DeOP [Han et al.(2023)Han, Zhong, Li, Han, and Ma]       80.8       88.2       74.6       38.2       38.0       38.4
      ZegCLIP [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu]       84.3       91.9       77.8       40.8       40.2       41.4
      OTSeg [Ye et al.(2024)Ye, Oh, et al.]       84.5       92.1       78.1       41.4       41.4       41.4
      Ours       ResNet101 [He et al.(2016)He, Zhang, Ren, and Sun]       85.3       87.7       83.1       42.5       42.6       42.4
Table 1: Comparison with others. Bold and underline indicates the best and the second-best.
Method hIoU sIoU uIoU
Baseline 24.6 31.8 20.0
+ SM 33.3 36.4 30.8
+ SM + MFE 36.3 36.8 35.8
+ SM + MFE + RQ 36.6 36.8 36.4
Table 2: Ablation on the proposed module.
Method hIoU sIoU uIoU
$\mathbf{F}_{u}$ 33.8 35.8 32.1
$\mathbf{F}_{u}$ + MLP 30.4 36.5 26.1
$\mathbf{C}_{u}$ 36.6 36.8 36.4
Table 3: Ablation on the candidate class embedding.
Structure of MFE hIoU sIoU uIoU
No Norm 35.8 36.2 35.3
BN 36.0 36.5 35.6
GN 36.6 36.8 36.4
Table 4: Ablation on the structure of MFE.
Implementation Details. The CLIP model applied in our method is based on the ViT-B/16 model, and the channel of the output text features is 512. All the experiments are conducted on 8 V100 GPUs, and the batch size is set to 16 for both datasets. The iterations are set to 20K and 80K for PASCAL VOC and COCO-Stuff. $\mathcal{L}_{cls}$ in $\mathcal{L}_{match}$ is focal loss [Lin et al.(2017)Lin, Goyal, Girshick, He, and Dollár] and the $\mathcal{L}_{mask}$ is a combination of IoU loss and DICE loss [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] in $\mathcal{L}_{match}$. We choose Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] with ResNet101 as the backbone with all other hyperparameters unchanged. 50 unseen and 50 random queries are added during inference. We apply the harmonic mean IoU (hIoU) following previous works [Zhou et al.(2023)Zhou, Lei, Zhang, Liu, and Liu] where $hIoU=\frac{2\cdot sIoU\cdot uIoU}{sIoU+uIoU}.$ as the metric where $sIoU$ and $uIoU$ indicate the mIoU of the seen classes and unseen classes, respectively. More details are in the Supplementary Materials.
4.1 Comparison with State-of-the-art methods
Table 1 presents an ablation study to evaluate the effectiveness of each proposed module: Split Matching (SM), Multi-scale Feature Enhancement (MFE), and the inclusion of additional randomly initialized queries. Starting from the baseline, the introduction of SM leads to a substantial improvement, increasing hIoU from 24.6 to 33.3 and uIoU from 20.0 to 30.8. Adding MFE further enhances the performance across all metrics, especially unseen classes (uIoU: 35.8). Finally, incorporating extra queries results in the best overall performance, achieving the highest scores in hIoU (36.6), sIoU (36.8), and uIoU (36.4). Note that our method achieves better performance even using a weaker backbone (ResNet101 vs. ViT-B).
Figure 5: Candidate query predictions visualization, with each row displaying images, GT, overall and candidate query predictions.
Figure 6: Visualization on predictions where each line shows the image, ground truth, ZegCLIP’s prediction, and ours.
4.2 Ablation Studies
To evaluate the effectiveness of our method, we conduct ablation studies on COCO-Stuff for 40K iterations using ResNet-50 as the backbone, with all hyperparameters unchanged. Additional experiments are provided in the Supplementary Materials due to space limitations.
Ablations on Proposed Modules. Table 4 summarizes the contributions of each proposed module. The baseline achieves suboptimal performance, with lower unseen IoU (uIoU) indicating limited generalization. Adding the Split Matching (SM) module significantly improves the model’s ability to capture unseen classes, as reflected in higher uIoU. The Multi-scale Feature Enhancement (MFE) further boosts the model’s performance by enhancing the interaction between queries and features. Finally, the inclusion of random queries leads to the best results across all metrics, demonstrating our contribution.
Ablations on Unseen Class Embedding. Table 4 presents an ablation study on the design of unseen class embeddings $\mathbf{C}_{u}$. Using raw dense features $\mathbf{F}_{u}$ from the backbone provides a reasonable baseline (uIoU: 32.1), while adding an MLP projection degrades performance, likely due to semantic distortion. In contrast, using CLS tokens from pseudo-masked regions yields the best results (uIoU: 36.4), suggesting that region-level CLS tokens better preserve semantic alignment. This also enables direct concatenation with seen-class text embeddings, maintaining cross-modal consistency without modality mismatch.
Ablations on structure of MFE. Table 4 presents the ablation study on the structure of the MFE module by comparing different normalization strategies: No Normalization (No Norm), Batch Normalization (BN), and Group Normalization (GN). Among the three, GN achieves the best performance across all metrics, yielding the highest hIoU (36.6), sIoU (36.8), and uIoU (36.4). These results demonstrate that incorporating Group Normalization into MFE significantly improves both seen and unseen segmentation performance.
4.3 Qualitative Analysis
Due to the space limitations, more results are in Supplementary Materials.
Visualization of candidate queries. Fig. 6 visualizes the predictions of candidate queries. Notably, these queries successfully activate on previously unannotated regions, enabling the model to localize unseen classes such as carrot and cow. This demonstrates that candidate queries can effectively discover latent categories and assign semantically correct class labels, even without explicit supervision.
Prediction Visualzation. Each row in Fig. 6 shows the input image, ground truth, ZegCLIP’s prediction, and ours. Our method successfully segments unseen categories such as “clouds,” “bushes,” and “playingfield,” which are missed or mislabeled by ZegCLIP. Notably, in both examples, the unseen class “clouds” is correctly identified by our model, demonstrating better generalization to unseen concepts.
5 Conclusion
In this work, we address the challenge of zero-shot semantic segmentation in query-based frameworks, where conventional Hungarian matching fails to account for unannotated objects. We propose Split Matching (SM), a novel mechanism that identifies and assigns unseen candidates from unannotated regions via CLIP-guided pseudo masks and class embeddings. Furthermore, we introduce an auxiliary Multi-scale Feature Enhancement (MFE) module to enhance the quality of key-value representations in the transformer decoder. Our method is the first to extend hungarian matching to the zero-shot setting and achieves state-of-the-art performance on two zero-shot segmentation datasets.